# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

### v1

@@ Abstract
In Part II, our optimized clustering (DPMM Cluster 2) yields a 14.04% detection rate (16 confirmed transits among 114 candidates) from the first five TESS SPOC sectors. This reflects a substantial enrichment over baseline rates: 0.4569% for the full TESS-SPOC project candidate set (7658 candidates across 1.68 million light curves), and 0.2650% for the FFI-based SPOC sample (7658 candidates across 2.89 million light curves; Doyle et al. 2024).
All computations were performed on a personal, CPU-only desktop with an Intel® Core™ i3-8100 processor and 32 GB RAM, using parallelized scoring and classification routines across four physical cores (see Section 5). CLARA processed over 87,000 TESS SPOC light curves (Sectors 1–5) without GPU acceleration. The complete codebase is available on Github.
keywords:
software: machine learning – exoplanets – methods: data analysis – astrometry – stars: general – (stars:) binaries (including multiple): close – planet–star interactions
pubyear: 2025pagerange: CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light CurvesD
@@ 1 Introduction
Recent advances in unsupervised machine learning for astronomical time-series analysis have shown promising results. Nun et al. (2015) demonstrated the effectiveness of Random Forest-based approaches for stellar variability classification, while Malanchev et al. (2021) explored deep learning methods for supernova discovery in large surveys. Active learning approaches have also emerged as powerful tools for personalized anomaly detection, with Lochner & Bassett (2016) introducing frameworks that combine human intuition with machine learning efficiency. The application of Unsupervised Random Forests (URFs) to exoplanet detection was demonstrated by Crake & Martínez-Galarza (2023) (MG23 hereafter), who showed that synthetic training sets could guide anomaly detection behavior in TESS light curves, with an approach optimizing models for the sharpest anomaly score tail. From MG23:
@@
However, these approaches have lacked systematic frameworks for controlling the discovery process and interpreting the physical nature of detected anomalies.
The challenge of anomaly detection in astrophysical time-series data is compounded by several factors: the high dimensionality of light curve features, the presence of various noise sources and systematic effects, and the need to distinguish between different classes of astrophysical phenomena Ivezić et al. (2019). Traditional clustering and dimensionality reduction techniques often fail to provide physically meaningful groupings of astronomical objects Ivezić et al. (2019), while purely data-driven approaches may discover statistically significant but astrophysically uninteresting patterns. Moreover, the analysis of time series photometry requires Lomb-Scargle periodogram techniques VanderPlas (2018). to properly characterize the temporal variability of astronomical sources that have been unevenly sampled.
We present CLARA (Controllable Learning for Anomaly Recognition in Astrophysics), a comprehensive pipeline that addresses these limitations through systematic synthetic set design and feature-to-metric mapping MG23. Our approach introduces novel methodologies for steering URF behavior through controlled variation of synthetic training parameters, enabling the targeted discovery of specific astrophysical phenomena. Furthermore, we develop morphological classification techniques using cosine similarity matching with known objects, and integrate astrometric data from Gaia DR3 Gaia Collaboration et al. (2023) to provide physical interpretation of discovered anomalies. This work aims to transform unsupervised anomaly detection from a purely statistical exercise into a physically guided discovery tool for astronomical surveys.
2 Guide for the Reader
@@
Part I (Section 3) investigates how synthetic light curve construction influences the performance of Unsupervised Random Forest (URF) models for anomaly detection in TESS 2-minute cadence SPOC light curves. It introduces a parameterized scoring metric controlled by a variable $\alpha$ that allows users to balance detection recall and astrophysical significance. We perform a grid search over synthetic set parameters (e.g., duration, noise level, cadence) and evaluate model behavior using TOI Recall, TOI Importance AUC, and Anomaly Rate. Generalization across sectors and benchmarking against prior URF variants (URF-1 to URF-3) are also presented.
@@
Part II (Section 4 onward) validates the astrophysical significance of high-scoring URF anomalies. We introduce a revised scoring scheme (Normalized Weighted Root Sum of Squares, or N-WRSS), and correlate anomaly scores with known stellar parameters (e.g., $T_{\mathrm{eff}}$, $\log g$, RUWE, $v_{\mathrm{tan}}$) from Gaia DR3. Morphological classification is performed using cosine similarity to labeled TOIs with SIMBAD object types (e.g., Pl, EB*, V*, PM*), allowing us to quantify the physical plausibility of detected signals.
Readers primarily interested in the machine learning framework and model tuning may focus on Part I, while those focused on astrophysical validation and interpretation may proceed directly to Part II. CLARA’s modular pipeline and evaluation methods are designed for both domains, with figures, tables, and code references included throughout to support reproducibility. For implementation details, datasets, and code repositories, please refer to the Data Availability section and the relevant Appendices.
Definitions and Terminology
We define the principal terms, abbreviations, and concepts used throughout the paper.
TESS: The Transiting Exoplanet Survey Satellite, a NASA mission that collects light curves from thousands of stars across multiple observing sectors, enabling large-scale time-domain photometry for exoplanet detection.
SPOC: Science Processing Operations Center. SPOC light curves are 2-minute cadence flux time series extracted and calibrated by NASA’s official TESS pipeline.
FFI: Full Frame Images. FFIs are 30-minute (later 10-minute and 200s) cadence image frames used to extract light curves across the entire TESS field of view. These are less targeted and more inclusive than SPOC.
TOI: TESS Object of Interest. A candidate planetary transit signal identified by the TESS pipeline and subjected to vetting.
URF: Unsupervised Random Forest. A machine learning model trained to distinguish real data from synthetic noise by learning an embedding space where outliers—defined by anomaly scores—can be identified.
Synthetic Set: A generated dataset of artificial light curves (typically noise-only or transit-injected) used to train the URF. The synthetic set acts as a pseudo-negative class for contrastive learning.
Anomaly Score: A numerical score produced by the URF that reflects how different a given light curve is from the synthetic training distribution. High scores indicate potential anomalies (e.g., transits, eclipses, instrumental artifacts).
WRSS: Weighted Root Sum Score. A statistical comparison metric used to evaluate score separability between models or datasets (e.g., between p3 and p9 URF variants).
Cosine Similarity Classification: A technique to match anomalous light curves against labeled reference examples (e.g., known eclipsing binaries or transits), using dimensionality-reduced feature vectors and cosine distance.
DTW: Dynamic Time Warping. A curve alignment method used to compare phase-folded light curves with variable durations and shifts.
Phase Folding: The process of folding a time series over a detected periodicity to align repeating patterns (e.g., planetary transits), enhancing signal-to-noise and making periodic features easier to identify.
DBSCAN: Density-Based Spatial Clustering of Applications with Noise. A clustering algorithm used to group similar anomalies and isolate dense clusters (e.g., eclipsing binaries, transit candidates).
DPMM: Dirichlet Process Mixture Model. A Bayesian non-parametric clustering technique that infers the number of clusters from the data and is used here to cluster anomaly scores or features.
Cluster -1: A reserved cluster label for outliers or noise points rejected by DBSCAN, typically enriched with rare or unique light curve morphology.
Morphological Similarity: The degree to which two light curves resemble each other in shape after normalization and alignment, used to propagate labels or identify classes.
Stratified Random Sampling: A sampling technique that divides the dataset into distinct subgroups (strata) based on a specific variable (e.g., anomaly score, sector, class label), and then samples proportionally from each group. This ensures balanced representation across the full range of a key feature or label, reducing sampling bias and preserving distributional structure.
[…]
@@
3 Exploring how the design of synthetic sets affects unsupervised anomaly detection
@@
Unsupervised random forest models for anomaly detection were first used in astronomical datasets in Baron & Poznanski (2017) where isolation forests were applied to SDSS galaxy data to identify the most unusual and potentially rare objects in the survey. As demonstrated ahead, the controllability of unsupervised anomaly detection hinges on our ability to systematically design synthetic training sets that guide model behavior toward discovering specific types of astrophysical phenomena. By varying key parameters such as the number of synthetic curves in the artificially generated data set, duration of curves/transits, frequency, and noise characteristics in our synthetic light curves, we can steer Unsupervised Random Forests to prioritize different anomaly signatures and effectively "tune" the discovery process for targeted scientific objectives.
3.1 Feature of a Curve
@@
The characteristic of a single light curve for input to our URF (Unsupervised Random Forest) model experiments follows established methods from our foundational paper MG23 where the first n=3000 flux points are stacked on top of nls = 1000 Lomb-Scargle periodogram, VanderPlas (2018) frequency points calculated using the LombScargle function of the astropy.timeseries module. The feature is thus an array with 4000 values. This function calculates the power of a time period where the power describes the extent of the periodicity of the selected time value, given by:
$P_{LS}(\omega)=\frac{1}{2}\left[\frac{\left(\sum_{j}(x_{j}-\bar{x})\cos(\omega(t_{j}-\tau))\right)^{2}}{\sum_{j}\cos^{2}(\omega(t_{j}-\tau))}\right.\\ \left.+\frac{\left(\sum_{j}(x_{j}-\bar{x})\sin(\omega(t_{j}-\tau))\right)^{2}}{\sum_{j}\sin^{2}(\omega(t_{j}-\tau))}\right]$ (1)
where:
$P_{LS}(\omega)$ $=\text{Lomb-Scargle power at angular frequency }\omega$
$x_{j}$ $=\text{observed flux values at times }t_{j}$
$\bar{x}$ $=\text{mean of the observed flux values}$
$\tau$ $=\text{time offset defined by:}$
$\tan(2\omega\tau)$ $=\frac{\sum_{j}\sin(2\omega t_{j})}{\sum_{j}\cos(2\omega t_{j})}$
An interesting property of the Lomb-Scargle periodogram is that it can, in principle, fail to detect strictly periodic signals if the sampling intervals are strictly irrational. However, this is more of a mathematical curiosity than a practical concern. In real-world observations, sampling intervals are always rational (due to finite precision) and periodic signals, even if originally irrational in frequency, are effectively rationalized by floating-point limitations in computation. Additionally, computing a Nyquist frequency becomes ill-defined under irrational sampling intervals, further complicating standard aliasing interpretations - but again, this rarely arises in empirical datasets.
@@ 3.2 URF Hyperparamter Randomized Search via Real + Synthetic Curve Stacking
For the real set (real features class), we took features of 1500 real curves from sector 2. A RandomizedSearchCV is performed over a predefined hyperparameter space for the RandomForestClassifier. The search is:
  • Randomized (n_iter=30)
  • Parallel (n_jobs=-1)
  • Reproducible (random_state=42)
The hyper-parameter search space is mentioned in Appendix A.2
@@
a. URF1. We generated the synthetic set by uniformly sampling feature values within the range of the real dataset. This establishes a baseline benchmark method.
X_fake = np.random.uniform(np.min(X_real), np.max(X_real), size=X_real.shape)
b. URF2. We used fixed hyper parameters mentioned in MG23. The following hyper parameters were used:
  • n_estimators = 100
  • max_features = "sqrt"
  • max_depth = 700
  • min_samples_split = 4
  • min_samples_leaf = 2
  • bootstrap = False
  • warm_start = True
c. URF3. We used a TOI based synthetic class where we obtained feature vectors from Tess Objects of Interest (TOIs) from Sector 1 of SPOC data, and set those features as the synthetic set for hyper-parameter search.
d. URF4. The synthetic set of features were computed from a set of synthetic light curves generated using transit models (box-shaped model and Mandel-Agol) injected into realistic noise baselines. The computational implementation of this was performed using the batman pypi package.
The four URF variants were designed to systematically explore different approaches to synthetic set construction for transit discovery. URF1 serves as a baseline control, using uniform random sampling within the feature space bounds to establish minimum performance expectations. URF2 tests the generalizability of the hyper parameters established in MG23, evaluating whether their optimization for stellar variability classification translates effectively to transit detection in TESS data. URF3 explores a realistic "ground truth" approach by using feature vectors from confirmed TOIs as the synthetic class, providing insight into the performance ceiling when the synthetic set closely matches actual transit signatures. Finally, URF4 introduces our controllable framework, where synthetic light curves are generated using physically motivated transit models with systematically variable parameters (ratio of real curves - 1500 - to synthetic curves - 300), duration, cadence, and noise characteristics), enabling targeted optimization for specific discovery objectives and forming the foundation for our systematic parameter exploration. For URF4, we fixed the real set to 1500 TESS SPOC curves across all experiments; investigating how the size and composition of this real subset affects model performance is a valuable direction for future work.
@@ 3.3 Anomaly Scoring
Following Baron & Poznanski (2017) and MG23, our URF implementation uses terminal node population as the anomaly scoring heuristic. Each tree in the forest is trained to distinguish real light curves from synthetic ones, with anomaly scores computed by tracking the population composition of terminal nodes. For each real light curve, we calculate a similarity score S as the average fraction of real data points sharing the same terminal node across all trees in the forest. The final anomaly score is defined as 1 - S, where a score of 1 indicates maximum anomaly (the object consistently lands alone in terminal nodes) and 0 indicates minimal anomaly (the object is always grouped with the majority of real data). This population-based approach provides a natural measure of how isolated each light curve appears relative to the broader dataset distribution.
URF Variant TP FP TP+FP TN Precision (%) Recall (%)
URF-1 22 2159 2181 0 1.01 12.57
URF-2 45 3990 4035 0 1.12 25.71
URF-3 14 300 314 15499 4.46 7.18
URF-4 64 4982 5046 0 1.27 36.57
3.4 Initial Results
Table 1 summarizes the performance of our four URF variants on TOI detection across the test dataset. URF-1, URF-2, and URF-4 demonstrate similar scoring behavior, assigning zero scores to most light curves and treating only non-zero scores as anomalies. URF-1 and URF-2 achieve baseline recalls of 12.6% and 25.7% respectively with 1̃% precision, while URF-4 shows improved performance at 36.6% recall and 1.3% precision. URF-3 exhibits fundamentally different behavior, assigning non-zero scores to nearly all light curves, requiring threshold-based evaluation (scores > 0.22, chosen arbitrarily based on figure 3) to define meaningful anomaly regions. When evaluated under these conditions, URF-3 achieves 7.2% recall with 4.5% precision, though this performance is best assessed on independent sectors to avoid overfitting bias from using TOI features as synthetic training data. The distinct scoring profiles highlight URF-4’s advantage in providing both strong recall performance and systematic controllability without the threshold sensitivity issues observed in URF-3
Figure 2: URF3 TOI (top) ranging from 0.0 to 0.65 with a normal shaped distribution ($\mu\approx$ 0.45) and All Curves (bottom) Anomaly Score with a normal shaped distribution ($mu\approx 0.1$) for Sector 1 Test Set showing distinct background separation between TOI anomaly scores and non TOI anomaly scores, interpreted as a result of URF overfitting to sector 1 TOI features as proved by same distribution plotted for sector 2 test set as in Figure 3
Figure 3: URF3 TOI (top) normal shaped distribution ($\mu\approx 0.1$) and All Curves (bottom) Anomaly Score normal shaped distribution ($\mu\approx 0.1$) for Sector 2 Test Set, showing TOI distribution (top) with near perfect distribution value and shape overlap over the entire anomaly score distribution (bottom)
URF-3 showed significant overfitting to TOI curve features of sector 1 which resulted in the distinct separation of TOIs from normal curve background when URF3 scored on sector 1. This was proved by cross sector evaluation of URF3 on SPOC sector 2 curve set. Refer to appendix A.1

### v3

@@ Abstract
In Part II, our optimized clustering (DPMM Cluster 2) yields a 14.04% detection rate (16 confirmed transits among 114 candidates) from the first five TESS SPOC sectors. This reflects a substantial enrichment over baseline rates: 0.4569% for the full TESS-SPOC project candidate set (7658 candidates across 1.68 million light curves), and 0.2650% for the FFI-based SPOC sample (7658 candidates across 2.89 million light curves;). Additionally, we perform Monte Carlo injection-recovery tests to assess feasibility, stability, justification and quantification of methods outlined.
All computations were performed on a personal, CPU-only desktop with an Intel® Core™ i3-8100 processor and 32 GB RAM, using parallelized scoring and classification routines across four physical cores (see Appendix A). CLARA processed over 87,000 TESS SPOC light curves (Sectors 1–5) without GPU acceleration. The complete codebase is available on Github.
[…]
software: astropy, lightkurve, batman, scikit-learn
@@ 1 Introduction
Recent advances in unsupervised machine learning for astronomical time-series analysis have shown promising results. I. Nun et al. (2015) demonstrated the effectiveness of Random Forest-based approaches for stellar variability classification, while K. L. Malanchev et al. (2021) explored deep learning methods for supernova discovery in large surveys. Active learning approaches have also emerged as powerful tools for personalized anomaly detection, with M. Lochner & B. A. Bassett (2016) introducing frameworks that combine human intuition with machine learning efficiency. The application of Unsupervised Random Forests (URFs) to exoplanet detection was demonstrated by D. A. Crake & J. R. Martínez-Galarza (2023) (MG23 hereafter), who showed that synthetic training sets could guide anomaly detection behavior in TESS light curves, with an approach optimizing models for the sharpest anomaly score tail. From MG23:
@@
However, these approaches have lacked systematic frameworks for controlling the anomaly discovery process and interpreting the physical nature of detected anomalies.
The challenge of anomaly detection in astrophysical time-series data is compounded by several factors: the high dimensionality of light curve features, the presence of various noise sources and systematic effects, and the need to distinguish between different classes of astrophysical phenomena Ž. Ivezić et al. (2019). Traditional clustering and dimensionality reduction techniques often fail to provide physically meaningful groupings of astronomical objects, while purely data-driven approaches may discover statistically significant but astrophysically uninteresting patterns and often need large sets of training data for targetted discovery. Then there are physics driven algorithms like Box Least Squares and Transit-Least-Squares which provide fixed detection strategies.
We present CLARA (Controllable Learning for Anomaly Recognition in Astrophysics), a comprehensive unsupervised random forest model design pipeline that addresses these limitations through systematic synthetic set design and design-feature-to-performance-metric mapping. Our approach introduces novel methodologies for steering URF behavior through controlled variation of synthetic set generation parameters, enabling the targeted discovery of specific classes of astrophysical phenomena characterized by light curve morphology. This is an unsupervised anomaly discovery algorithm with controllable behavior and semi-supervised downstream filtering for astrophysical contextualization of anomalies found. Furthermore, we develop morphological classification techniques using cosine similarity matching with known objects and integrate astrometric data from Gaia DR3 Gaia Collaboration et al. (2023) to provide physical interpretation of discovered anomalies. This work aims to transform unsupervised anomaly detection from a purely statistical exercise into a physically guided discovery and prioritization tool for astronomical surveys.
@@
Part I investigates how synthetic light curve construction for fine tuning of hyper-parameters influences the performance of Unsupervised Random Forest (URF) models for anomaly detection in TESS 2-minute cadence SPOC light curves. It introduces a parameterized scoring metric controlled by a variable coefficient $\alpha$ that allows balancing of detection recall and astrophysical significance. We perform a grid search over synthetic set parameters (e.g., duration, noise level, cadence) and predict model behavior using TOI (Tess Object of Interest) Recall, Importance metrics for TOIs recovered by URF models, Anomaly Rate and other performance metrics over a small stratified test set. Then we establish generalization of the predicted model performance across both similar subsets from the same sector and diverse populations across five different sectors, presented later.
@@
Part II validates the astrophysical significance of high-scoring URF anomalies. We introduce a revised scoring scheme (Normalized Weighted Root Sum of Squares, or N-WRSS), and correlate anomaly scores with known stellar parameters (e.g., $T_{\mathrm{eff}}$, $\log g$, RUWE, $v_{\mathrm{tan}}$) from Gaia DR3. Morphological classification is performed using cosine similarity to small representative set of labeled TOI features with SIMBAD object types (e.g., Pl, EB*, V*, PM*), allowing us to quantify the physical plausibility of detected signals. Then we finish by introducing an advance 10 dimensional feature set for enhanced clustering of transit-like curves to filter out non-transit-like curves.
A key point to note is that rather than competing with established transit detection methods, CLARA demonstrates how unsupervised ML frameworks can be systematically tuned and validated for specific discovery objectives using synthetically generated data - a capability with broad applications across astronomical time-domain surveys, including prioritization of targets when faced with limited follow-up resources, identification of rare or anomalous events in large datasets that have high probability of fitting an astrophysical class, and adaptive survey strategies that can evolve based on emerging scientific priorities or unexpected discoveries.
2 Methods
@@
2.1 Part 1 - Exploring how the design of synthetic sets affects unsupervised anomaly detection
@@
Unsupervised random forest models for anomaly detection were first used in astronomical datasets in D. Baron & D. Poznanski (2017) where isolation forests were applied to SDSS galaxy data to identify the most unusual and potentially rare objects in the survey. As demonstrated ahead, the controllability of unsupervised anomaly detection hinges on our ability to systematically design synthetic training sets that guide model behavior toward discovering specific types of astrophysical phenomena. By varying key parameters such as the number of synthetic curves in the artificially generated data set, duration of curves/transits, frequency, and noise characteristics in our synthetic light curves, we can predict the generalizable behavior of an URF model by testing on a small labeled representative dataset, allowing us to essentially steer Unsupervised Random Forests to prioritize different anomaly signatures and effectively ”tune” the discovery process for targeted scientific objectives ranging from higher detection rate to better scoring of anomalies.
2.2 Feature of a Curve - I
@@
The characteristic of a single light curve for input to our URF (Unsupervised Random Forest) model experiments follows established methods from our foundational paper D. A. Crake & J. R. Martínez-Galarza (2023) where the first $n_{flux}$=3000 flux points are stacked on top of $n_{power}$=1000 Lomb-Scargle periodogram, J. T. VanderPlas (2018) frequency points calculated using the LombScargle function of the astropy.timeseries module. The Lomb-Scargle periodogram is a method for detecting and estimating the strength of periodic signals in unevenly sampled time series data, extending the classical Fourier analysis to irregularly spaced observations as characteristic of TESS light curve data. This function calculates the power of a period where the power describes the extent of the periodicity of the selected time value. Hence the feature is essentially a 4000 dimensional vector.
@@ 2.3 URF Hyperparamter Randomized Search via Real + Synthetic Curve Stacking
For the real set (real features class), we took features (Section 2.2) of 1500 real curves from TESS SPOC sector 2. A RandomizedSearchCV is performed over a predefined hyperparameter space for the RandomForestClassifier. The hyper-parameter search space is defined in Appendix B.2
@@
a. URF1. We generated the synthetic set by uniformly sampling feature values within the range of the real dataset as a baseline benchmark.
[…]
b. URF2. We used fixed hyperparameters from MG23 with n_estimators=100, max_features=”sqrt”, max_depth=700, min_samples_split=4, min_samples_leaf=2, bootstrap=False, and warm_start=True.
c. URF3. We used feature vectors from confirmed Tess Objects of Interest (TOIs) from Sector 1 of SPOC data as the synthetic set for hyperparameter search.
d. URF4. The features were computed from a set of synthetic light curves generated using transit models (box-shaped model and Mandel-Agol) injected into realistic noise baselines. The computational implementation of this was performed using the batman pypi package.
The four URF variants were designed to systematically explore different approaches to synthetic set construction for transit discovery. URF1 serves as a baseline control, using uniform random sampling within the feature space bounds to establish minimum performance expectations. URF2 tests the generalizability of the hyperparameters established in MG23, evaluating whether their optimization for stellar variability classification translates effectively to transit detection in TESS data. URF3 explores a realistic ”ground truth” approach by using feature vectors from confirmed TOIs as the synthetic class, providing insight into the performance ceiling when the synthetic set closely matches actual transit signatures. Finally, URF4, our main URF variant for further work, was chosen because of its inherent controllability over physically motivated synthetic dataset design, where synthetic light curves are generated using transit models with systematically variable parameters (ratio of real curves - 1500 - to synthetic curves - 300), duration, cadence, and noise characteristics), enabling targeted optimization for specific discovery objectives and forming the foundation for our systematic parameter exploration. For URF4, we fixed the real set to 1500 TESS SPOC curves across all experiments; investigating how the size and composition of this real subset affects model performance is a valuable direction for future work. URF 1-4 model performance is as listed in Table 10.
@@ 2.4 Anomaly Scoring
Following D. Baron & D. Poznanski (2017) and D. A. Crake & J. R. Martínez-Galarza (2023), our URF implementation uses terminal node population as the anomaly scoring heuristic. Each tree in the forest is trained to distinguish real light curves from synthetic ones, with anomaly scores computed by tracking the population composition of terminal nodes. For each real light curve, we calculate a similarity score S as the average fraction of real data points sharing the same terminal node across all trees in the forest. The final anomaly score is defined as 1 - S, where a score of 1 indicates maximum anomaly (the object consistently lands alone in terminal nodes) and 0 indicates minimal anomaly (the object is always grouped with the majority of real data). This population-based approach provides a natural measure of how isolated each light curve appears relative to the broader dataset distribution.